OpenStudy (anonymous):

I just want someone to double check my answer And i will give medals what is the graph of the function y = -2 * 5^x

OpenStudy (anonymous):

You are right. At y=-10, x is 1. -10=-2*5^1 -10=-2*5 -10=-10

OpenStudy (anonymous):

I also looked at the graph, and it is